OpenStudy (anonymous):

Find dy/dx for y = ∫(1+ sin^2u)/(1 + cos^2u) interval: [π, π + x]

OpenStudy (anonymous):

Fundamental Theorem of Calculus problem

OpenStudy (science0229):

Let \[g(x)=\frac{ 1+\sin^2x }{ 1+\cos^2x }\] and define G(x) in which\[G'(x)=g(x)\]

OpenStudy (science0229):

Then the problem becomes\[\int\limits_{x}^{\pi+x}g(u)du=f(x)\]

OpenStudy (science0229):

Using our definition of G(x),\[\int\limits_{x}^{\pi+x}g(u)du=G(\pi+x)-G(x)=f(x)\]

OpenStudy (science0229):

We have to find f'(x), so...\[f'(x)=(G(\pi+x)-G(x))'=g(\pi+x)-g(x)\]

OpenStudy (science0229):

You can do the rest since we defined\[g(x)=\frac{ 1+\sin^2x }{ 1+\cos^2x }\]

OpenStudy (science0229):

From this problem, we can derive the formula that if\[f(x)=\int\limits_{a}^{b}g(x)dx\] then \[f'(x)=g(b)-g(a)\]